# Deep Neural Crossover

### v1

@@ 4.1. Domains
[…]
@@
Instance
Optimal Solution
DNC DNC-PT One Point
Adaptive Uniform
Equiprobable Uniform
DNC-MP Multi-Parent
games120
9
13.25 19.5 34.45
33.2
33.65
12.0 28.7
myciel7
8
25.5 27.4 64.75
63.0
63.9
18.7 55.2
miles1000
42
47.9 48.0 55.1
54.0
57.45
47.16 55.2
miles1500
73
76.55 77.0 78.85
77.65
82.0
76.2 76.85
mulsol.i.2.col
31
41.55 39.0 71.2
68.3
72.35
34.75 61.2
queen8_12
12
19.15 19.5 31.4
30.8
31.8
18.8 28.5
zeroin.i.1.col
49
56.65 - 80.65
79.15
83.25
52.46 71.75
zeroin.i.2.col
30
40.8 39.0 77.15
75.8
77.0
37.6 67.05
@@
Instance DNC DNC-PT One Point
Adaptive Uniform
Equiprobable Uniform
DNC-MP Multi-Parent
BPP_14 0.8 0.793 0.727
0.628
0.776
0.868 0.856
BPP_181 0.773 0.762 0.697
0.583
0.763
0.866 0.849
BPP_195 0.811 - 0.705
0.598
0.756
0.883 0.853
BPP_359 0.768 0.761 0.699
0.572
0.761
0.852 0.846
BPP_360 0.79 0.762 0.724
0.63
0.77
0.85 0.839
BPP_40 0.821 0.799 0.728
0.647
0.788
0.88 0.856
BPP_47 0.75 0.755 0.683
0.582
0.745
0.832 0.829
BPP_60 0.794 0.796 0.726
0.62
0.778
0.868 0.859
BPP_645 0.814 0.822 0.735
0.653
0.781
0.886 0.876
BPP_785 0.783 0.78 0.7
0.581
0.752
0.867 0.848
BPP_832 0.831 0.823 0.753
0.638
0.791
0.886 0.874

### v2

@@ 4.1. Domains
We employ an integer encoding scheme where the $i$-th gene corresponds to the $i$-th item, and each gene represents the bin number to be packed into. As before, we set a fitness score of $\infty$- ∞ to invalid solutions (i.e., solutions in which some bin is assigned with more weight than the specified capacity $C$). For valid solutions the fitness function is: $\frac{\sum_{i=1}^{N}(\frac{\textit{fill}_{i}}{C})^{2}}{N}$. Where $N$ denotes the number of bins used by the solution, $\textit{fill}_{i}$ denotes the sum of weights of the items in the $i$-th bin, and $C$ marks the bin capacity. Regarding valid solutions, the fitness score varies between 0 and 1, with bigger values being better solutions.
@@
Instance Optimal Solution DNC DNC-PT One Point Adaptive Uniform Equiprobable Uniform DNC-MP Multi-Parent
games120 9 13.25 19.5 34.45 33.2 33.65 12.0 28.7
myciel7 8 25.5 27.4 64.75 63.0 63.9 18.7 55.2
miles1000 42 47.9 48.0 55.1 54.0 57.45 47.16 55.2
miles1500 73 76.55 77.0 78.85 77.65 82.0 76.2 76.85
mulsol.i.2.col 31 41.55 39.0 71.2 68.3 72.35 34.75 61.2
queen8_12 12 19.15 19.5 31.4 30.8 31.8 18.8 28.5
zeroin.i.1.col 49 56.65 - 80.65 79.15 83.25 52.46 71.75
zeroin.i.2.col 30 40.8 39.0 77.15 75.8 77.0 37.6 67.05
@@
Instance DNC DNC-PT One Point Adaptive Uniform Equiprobable Uniform DNC-MP Multi-Parent
BPP_14 0.8 0.793 0.727 0.628 0.776 0.868 0.856
BPP_181 0.773 0.762 0.697 0.583 0.763 0.866 0.849
BPP_195 0.811 - 0.705 0.598 0.756 0.883 0.853
BPP_359 0.768 0.761 0.699 0.572 0.761 0.852 0.846
BPP_360 0.79 0.762 0.724 0.63 0.77 0.85 0.839
BPP_40 0.821 0.799 0.728 0.647 0.788 0.88 0.856
BPP_47 0.75 0.755 0.683 0.582 0.745 0.832 0.829
BPP_60 0.794 0.796 0.726 0.62 0.778 0.868 0.859
BPP_645 0.814 0.822 0.735 0.653 0.781 0.886 0.876
BPP_785 0.783 0.78 0.7 0.581 0.752 0.867 0.848
BPP_832 0.831 0.823 0.753 0.638 0.791 0.886 0.874